# From Noise to Laws: Regularized Time-Series Forecasting via Denoised Dynamic Graphs

### v1

@@ 5 Conclusion
Appendix A Acknowledgments
The authors used large language models solely for language polishing and grammar editing. All technical content, methods, experiments, and analysis were conducted entirely by the authors.
Appendix B Datasets
We evaluate DORIC on six real-world benchmarks, covering the five domains of energy, traffic, economics, weather, and disease. We use the same datasets as (Wu et al., 2021), and provide additional information in Table 4, as given in the original Autoformer paper.
Table 4: Descriptions of the datasets
@@ Appendix B Datasets
Appendix C Proofs for Propositions
@@
since $\|\bar{A}_{t}\|_{2}\leq\rho(\bar{A}_{t})\leq 1$ by normalization. Summing both contributions yields equation 11. The nonlinearity bound follows from the $1$-Lipschitz property of $\sigma$, and the product bound from the Lipschitz constant of compositions. ∎
@@
Combining Props. C2 yields a two-level stability picture: (i) Temporal contraction along the horizon due to the RD step whenever $\kappa+\gamma<1$ (uniformly over time, with a perturbation margin $\kappa\varepsilon<\gamma$ for graph noise); (ii) Spatial Lipschitz control within each window via explicit operator-norm constraints on $W_{\mathrm{self}}^{(\ell)},U_{\mathrm{nei}}^{(\ell)}$. In particular, if $\|W_{\mathrm{self}}^{(\ell)}\|_{2}+\|U_{\mathrm{nei}}^{(\ell)}\|_{2}<1$ for all $\ell$, the stacked graph encoder is a contraction on $(\mathbb{R}^{D\times d},\|\cdot\|_{2})$, complementing the temporal contraction of the RD transition and explaining stable, well-conditioned rollouts over long horizons under the loss terms of Eq. (9).
Appendix D Pseudo-code of PRISM
@@
Appendix E Further Ablation Studies
@@ E.1 Quantitative extensions
E.2 Mechanism-level diagnostics
@@
E.3 Explanatory figures (reproducible from body tables)
@@
Figure 3: Ablation deltas computed from the body tables (exact values reproduced).
@@
Figure 4: Average relative degradation across datasets (%); derived from body ablations.
E.4 Discussion: How ablations map to mechanisms
@@
Appendix F Adjacency Structure Analysis (Thresholded Correlations)
@@
What to read from the heatmaps.(Figure  5)
@@
F.1 Dataset-specific interpretations
@@
F.2 Consistency checks and failure modes
@@
F.3 What the matrices imply for forecasting
@@
Figure 5: Thresholded correlation adjacencies used by PRISM. Bright cells survive $|C_{t}|>\tau$ and are reweighted by $|C_{t}|^{\gamma}$; black cells are pruned. Self-loops are added only after normalization when forming $\bar{A}_{t}$.

### v2

@@ 5 Conclusion
Appendix A Datasets
We evaluate DORIC on six real-world benchmarks, covering the five domains of energy, traffic, economics, weather, and disease. We use the same datasets as (3), and provide additional information in Table A.1, as given in the original Autoformer paper.
Table A.1: Descriptions of the datasets
@@ Appendix A Datasets
Appendix B Proofs for Propositions
@@
since $\|\bar{A}_{t}\|_{2}\leq\rho(\bar{A}_{t})\leq 1$ by normalization. Summing both contributions yields equation B. The nonlinearity bound follows from the $1$-Lipschitz property of $\sigma$, and the product bound from the Lipschitz constant of compositions. ∎
@@
Combining Props. B2 yields a two-level stability picture: (i) Temporal contraction along the horizon due to the RD step whenever $\kappa+\gamma<1$ (uniformly over time, with a perturbation margin $\kappa\varepsilon<\gamma$ for graph noise); (ii) Spatial Lipschitz control within each window via explicit operator-norm constraints on $W_{\mathrm{self}}^{(\ell)},U_{\mathrm{nei}}^{(\ell)}$. In particular, if $\|W_{\mathrm{self}}^{(\ell)}\|_{2}+\|U_{\mathrm{nei}}^{(\ell)}\|_{2}<1$ for all $\ell$, the stacked graph encoder is a contraction on $(\mathbb{R}^{D\times d},\|\cdot\|_{2})$, complementing the temporal contraction of the RD transition and explaining stable, well-conditioned rollouts over long horizons under the loss terms of Eq. (9).
Appendix C Pseudo-code of PRISM
@@
Appendix D Further Ablation Studies
@@ D.1 Quantitative extensions
D.2 Mechanism-level diagnostics
@@
D.3 Explanatory figures (reproducible from body tables)
@@
Figure D.1: Ablation deltas computed from the body tables (exact values reproduced).
@@
Figure D.2: Average relative degradation across datasets (%); derived from body ablations.
D.4 Discussion: How ablations map to mechanisms
@@
Appendix E Adjacency Structure Analysis (Thresholded Correlations)
@@
What to read from the heatmaps.(Figure  E.3)
@@
E.1 Dataset-specific interpretations
@@
E.2 Consistency checks and failure modes
@@
E.3 What the matrices imply for forecasting
@@
Figure E.3: Thresholded correlation adjacencies used by PRISM. Bright cells survive $|C_{t}|>\tau$ and are reweighted by $|C_{t}|^{\gamma}$; black cells are pruned. Self-loops are added only after normalization when forming $\bar{A}_{t}$.